OpenStudy (anonymous):

if f(x)= 6x-9 what is f of f (-1)

OpenStudy (anonymous):

f(-1)=6(-1)-9 f(-1)=-6-9 so f(-1)=-15

OpenStudy (anonymous):

hey can you help me with some similar problems please

OpenStudy (anonymous):

yeah:)

OpenStudy (anonymous):

What is the problem Breno?

OpenStudy (anonymous):

if f(x)= 1/x what is f of f(x)

OpenStudy (anonymous):

Same thing, f(x)=6x-9 f(1/x)= 6(1/x)-9 f(1/x)= (6/1)(1/x)-9 f(1/x)= (6/x)-9 ^.^

OpenStudy (anonymous):

I think I confuse your question, lol, sorry... Ok, so if f(x)= 1/x what is f of f(x)= -x

OpenStudy (anonymous):

huh

OpenStudy (mathstudent55):

f(x) = 1/x f(f(x)) = f(1/x) = 1/(1/x) = x

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

Given that f(x)=|x| and g(x)=2x−3 what is f of g(x)

OpenStudy (mathstudent55):

f(g(x)) = f(2x - 3) = |2x - 3|

OpenStudy (anonymous):

i got it nvm

OpenStudy (mathstudent55):

You mean f = |x|, what is f(f(x))?